4 stion n an experiment on hydrolysis, the pH values of 0.1M sodium acetate (CHhcoONa) and 0.1 chiorid recorded using pH paper of differentranges. Use these values to determine the correct pH then com Table Abelow DATA AND CALOULATONNS Table A MOH1 Paper Range Wide Range Paper Narrow Range Paper 0.1 MCHvcooNa 4 60 90 95 11 12 Calculations Part B Insert the net ionic equatioms, equilibriumconstant expressions, K. and Ka for the hydrolysable ions, in Table B. Show calculations for K and Ka (see next Pagel. Table B Equilibrium Constant Expression Net Ionic Equation Value 0.1 M NH 12 points) pointsl
Answer #1

Part A:

CH3COONa is salt of weak acid strong base so its aqueous solution must be basic in nature. pH will be 9.5

NH4Cl is salt of strong acid weak base so its aqueous solution will be acidic in nature, its pH from above data is 4.5

Now [H+] for CH3COONa =

we have pH= -log10 [H+]

So, [H+]= antilog[-pH]=antilog(-9.5)= 3.16\times10-10

now [H+][OH-]=10-14 now [OH-]=10-14/3.16\times10-10 = 3.16\times10-5

For NH4Cl

So, [H+]= antilog[-pH]=antilog(-4.5)= 3.16\times10-5

now [H+][OH-]=10-14 now [OH-]=10-14/3.16\times10-5 = 3.16\times10-10